OpenStudy (anonymous):

The graph shows the number of sprays an automatic air freshener dispenses, y, in x minutes:A graph is shown. On the x-axis, the values are from 0 till 150 in increments of 30 for each grid line, and on the y-axis, the values are from 0 to 10 in increments of 2 for each grid line. The title of the graph on the x-axis is Time in minutes, and the title on the y-axis is Number of Sprays. A line is shown connecting ordered pairs 30, 2 and 60, 4 and 90, 6 and 120, 8. The title for the graph is Dispense Rate.Which expression can be used to calculate the rate per minute @which the airfreshener sprays?

OpenStudy (anonymous):

Possible answers below fraction 30 over 2 fraction 2 over 30 fraction 30 over 8 fraction 8 over 30

OpenStudy (johnweldon1993):

Well the rate of change between 2 points would be \[\large \frac{y_2 - y_1}{x_2 - x_1}\] So we'll pick 2 points...lets do (30,2) and (60,4) So we have \[\large \frac{60 - 30}{4 - 2}\] And simplify a little bit

OpenStudy (anonymous):

it increases by 30 every two minutes. The average rate of change is 30/2 This simplifies to 15/1

OpenStudy (anonymous):

So... the answer is 2 over 30..

OpenStudy (johnweldon1993):

Not quite...if you simplify my fraction above you will have \[\large \frac{30}{2}\]

OpenStudy (anonymous):

Okay , I've got it now :) Thanks for the help @johnweldon1993 and @bluediamond1999